Circular No. 6344

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) D. Schleicher, Lowell Observatory, reports: “Narrow-band photometry

of comet C/1996 B2 was obtained by R. Millis using the Hall 1.07-m telescope at Lowell Observatory on Mar. 19.4 UT (r = 1.16 AU). Haser-model production rates were log Q(OH) = 28.83, log Q(CN) = 26.35, and log Q(C_{2}) = 26.50. The dust-production rate was log (Afρ) = 3.72. Combining these results with those previously obtained at r = 1.90, 1.60, and 1.51 AU (cf. IAUC 6311, 6333) indicates that the OH production rate has finally begun to increase, with an r-dependent slope of −0.9 from 1.51 to 1.16 AU. During this same interval, the CN production had a slope of −1.7, after also being nearly flat. Both C_{2} and dust have exhibited steady increases since 1.90 AU, with a slope for C_{2} of −1.7 and a slope for Afρ = −2.5.”

L. Jorda, Meudon Observatory and Laboratoire d’Astronomie Spatiale, Marseille; J. Lecacheux, Meudon Observatory; and F. Colas, Bureau des Longitudes, Paris, write: “Subarcsecond imaging since Mar. 9 of the inner coma of this comet with the Pic-du-Midi 1.05-m telescope shows at least two curved jets rotating clockwise and changing orientation on a time scale of a few minutes. From a continuous broadband (500–900 nm) image sequence obtained during Mar. 18.97–19.23 under good seeing (0′′.9) conditions, the activity of 2000-km-long jets begins at the morning limb near p.a. 200^{o}, is at maximum at noon near p.a. 95^{o}, and disappears at the evening limb near p.a. 350^{o}. The rotation period seems to have been in the range 5–9 hr (and is possibly near 6.6 hr), from the comparison of our early images with the ESO-NTT image obtained on Mar. 19.342 (IAUC 6343). This suggests that the nucleus shows its southern hemisphere to earth-based observers and that the sun lies near its equatorial plane. Images are available at

URL http://www.bdl.fr/s2p/hyakutake.html.”

L. M. Woodney, University of Maryland (UM); J. McMullin, Submillimeter Telescope Observatory; and M. F. A’Hearn, UM, report the detection of OCS J(12-11) emission from the comet: “Observations were made of the OCS line on Mar. 19 (when Δ = 0.217 AU, with geocentric velocity −51.6 km/s) at the NRAO 12-m telescope on Kitt Peak (peak antenna temperature 15 ± 4 mK; area 0.0194 K km/s); the line had FWHM = 1.2 ± 0.4 km/s and was observed at −52.19 ± 0.16 km/s. Assuming kinetic temperatures of 10 and 30 K (and an expansion velocity of 0.8 km/s), we estimate the OCS production rate, log Q(OCS) = 26.4 and 25.7, respectively; comparing this with log Q(OH) = 28.83, measured by R. Millis on Mar. 19.4 (above), yields Q(OCS)/Q(OH) = 0.004 and 0.001, respectively.”

1996 March 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green